The sum of two numbers is 128. The greater number is 7 less than twice the smaller number. Find the numbers.
Answer by ReadingBoosters(3246)

x = greater number
y = smaller number
x + y = 128
x = 2y - 7
2y - 7 + y = 128
Combine like terms
3y - 7 = 128
3y - 7 + 7 = 128 + 7
3y = 135
3y/3 = 135/3
y = 45
x = 2(45)-7
x = 90-7
x = 83
Proof
83 + 45 = 128